50% pure 1 g CaCO3 is dissolved in 25 ml H2SO4 solution. A further 30 ml of 0.5 mol NaOH is required for complete dissolution of the mixture. So what was the concentration of H2SO4?

1
Expert's answer
2021-11-16T11:28:03-0500

CaCO3 + H2SO4 = H2CO3 + CaSO4

H2SO4 + 2NaOH = Na2SO4 + 2H2O


1) Active mass of CaCO3 is: 1 g x 0.5 = 0.5 g.

n (CaCO3) = m / M

M (CaCO3) = 100.1 g/mol

n (CaCO3) = 0.5 / 100.1 = 0.005 mol

According to the equation, n (H2SO4) = n (CaCO3).


2) CM = n / V

n (NaOH) = CM x V =0.5 x 0.03 = 0.015 mol

According to the equation, n (H2SO4) = 1/2 x n (NaOH) = 1/2 x 0.015 = 0.0075 mol.


As H2SO4 reacted with both CaCO3 and NaOH, the amount of used H2SO4 is:

n (H2SO4) = 0.005 + 0.0075 = 0.0125 mol


CM (H2SO4) = n / V = 0.0125 / 0.025 = 0.5 M